OpenStudy (anonymous):

Two urns each contain green balls and blue balls. Urn I contains 4 green balls and 6 blue balls, and Urn II contains 6 green balls and 2 blue balls. A ball is drawn at random from each urn. What is the probability that both balls are blue? A.2/51 B.3/20 C.1/10 D.4/153

OpenStudy (hoblos):

P= 6/10 * 2/8 = 12/80 =3/20